OpenStudy (anonymous):

what the slope in 4x+2y=6

OpenStudy (anonymous):

Well, get the y alone on the left side of the equation then divide out the number right next to it (the coefficient.) Do you know how to do that? O.o

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well how would you remove the 4x from the 4x+2y?

OpenStudy (unklerhaukus):

4x + 2y = 6 take away 4x form both sides of the equation then divide both sides of the equation by 2 you should now have the equation in the form y = mx + b where m is the slope, and b is the y-intercept

OpenStudy (anonymous):

May I suggest a shortcut?

OpenStudy (anonymous):

In \[Ax+Bx=C\] The slope is \[\frac{ -A }{ B }\] For example, if the equation was 6x-2y=4 You'd put 6 in for A, and -2 in for B. \[\frac{ -6 }{ -2 }=3\] So the slope of that equation is equal to 3. Just to check to make sure that this is true, we can rearrange the equation to do it in slope intercept form. \[6x-2y=4\] We want to get the -2y alone on one side, so we'd subtract the 6x to the other side. \[-2y=-6x+4\] Now, we want to get the y completely alone, without any numbers before it, so we'd divide by -2 on both sides. \[y=3x-2\] In both cases, the slope is 3. So \[\frac{ -A }{ B }\] Is a valid equation for slope from standard form.

OpenStudy (unklerhaukus):

*\[Ax+B\color{red}y = C \]

OpenStudy (anonymous):

So the question is 4x+2y=6 What you want to do is get y alone so subtract 4x from both sides and you get 2y=-4x+6 Next get y alone by dividing both sides by 2 2y/2=-4x+6/2 You then get y=-2x+3 The slope will always be the one with x so the slope is -2x (or -2x/1)

OpenStudy (unklerhaukus):

the slope does not depend on x

OpenStudy (anonymous):

Thanks unkle. c; I thought something looked off as I was writing it. But I think that Novalynne is not necessarily saying that the slope depends on x, just that it is paired with the x in slope-intercept form.

OpenStudy (anonymous):

Sorry for the confusion but that's exactly what I mean't Halorazer. Slope-intercept form is y=mx+b and slope is m so in this case the slope will always be with x